OpenStudy (ttp):

Parker earned $36 by selling 30 glasses of fresh-squeezed lemonade at his lemonade stand on a Saturday morning. He charged $1.00 for a small glass and $1.50 for a large glass. Use a model to determine how many large glasses he sold. 10 11 12 13

OpenStudy (the_fizicx99):

12 because 12*1.5 = 18 $1 per cup so 18 small cups = 18 + 18 = 36